OpenStudy (anonymous):

how do you find the center and radius of a circle with points (0,1) and (2,3).

OpenStudy (anonymous):

those points are on the circle

OpenStudy (anonymous):

(y-h)^2+(x-b)^2=r^2 r is radius (h,b) is center of circle

OpenStudy (anonymous):

but how can you use that if you don't already know the center or the radius

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lol im thinking ofeasiest way to do it

OpenStudy (anonymous):

haha okay, thank you

OpenStudy (anonymous):

1st equation: (1-h)^2+(0-b)^2=r^2 (1-h)^2+b^2=r^2 2nd equation: (3-h)^2+(2-b)^2=r^2 combine equations: (1-h)^2+b^2=(3-h)^2+(2-b)^2

OpenStudy (anonymous):

okay, thanks so much!!

OpenStudy (anonymous):

u still have to solve it tho lol

OpenStudy (anonymous):

haha yes, i know. but that helps a lot knowing how to find center

OpenStudy (anonymous):

after u solve that u should get: h+b=3 or some variant

OpenStudy (zarkon):

are (0,1) and (2,3) just two points on the circle?

OpenStudy (zarkon):

or is there something else we know

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you have only two point. you are going to need something else.

myininaya (myininaya):

not enough info i believe

OpenStudy (zarkon):

then the problem can't be solved

OpenStudy (anonymous):

all we know is that (0,1) and (2,3) are on the line. and we need to find center and radius. this is in my pre-calc textbook as a homework question

myininaya (myininaya):

i go to bed now i have to get up at 5:00 am

myininaya (myininaya):

line?

OpenStudy (anonymous):

good night

OpenStudy (anonymous):

oh a line. maybe we can do it then

OpenStudy (zarkon):

is it a diameter?

OpenStudy (anonymous):

myininaya! congratulations again. and you are so colorful!

myininaya (myininaya):

thanks :)

myininaya (myininaya):

you guys can do this one maybe if you guess correctly what he wants lol i will leave you

OpenStudy (anonymous):

@jahtoday post the entire problem and maybe we can solve it. two point and the line they are on is not enough

OpenStudy (anonymous):

its on a circle

OpenStudy (anonymous):

its not my post lol ive just been trying to solve it and i couldnt get any further

OpenStudy (anonymous):

The problem is to find the center and radius of the circle. Write the standard form of the equation. The graph has a circle with two points labeled and those two points are (2,3) and (0,1). The center is not labeled

OpenStudy (zarkon):

does the line go through the center?

OpenStudy (anonymous):

If you guys aren't busy can you help me with my question? Thanks :)

OpenStudy (anonymous):

there is no line, its a circle. however the two points are on the diameter of the line

OpenStudy (zarkon):

you mean diameter of the circle?

OpenStudy (anonymous):

yes, diameter of the circle.

OpenStudy (zarkon):

then the problem is simple

OpenStudy (anonymous):

*smacks head against wall*!

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

i just graphed it, and on my graph, because (0,1) and (2,3) was diameter, my graph shows (1,2) as center. is that correct?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

sorry, i should've explained more at the beginning :/

OpenStudy (anonymous):

and then i can plug into the equation to find the radius?

OpenStudy (zarkon):

what is the distance from the center to either of the points...

OpenStudy (anonymous):

i haven't figured that out yet, do i need to?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

what will i use that for?

OpenStudy (zarkon):

\[D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

but how will that help me find my radius?

OpenStudy (zarkon):

the radius of a circle is the distance from the center of the circle to any point on the circle

OpenStudy (anonymous):

oh so you can just divide by two?

OpenStudy (zarkon):

no...you need to use the distance formula

OpenStudy (anonymous):

once you use the distance formula, you divide by two right?

OpenStudy (zarkon):

\[r=\sqrt{(2-1)^2+(3-2)^2}=\sqrt{1^2+1^2}=\sqrt{1+1}=\sqrt{2}\]

OpenStudy (anonymous):

i also just plugged one of my coordinates and the center into the standard form and solved for radius, i believe it ends up being +/- square root of 2

OpenStudy (anonymous):

we ended up with same thing! thank you so much :)

OpenStudy (zarkon):

that is because the formula for the circle is really the same as the distance formula

OpenStudy (anonymous):

but i have one more question!! in the general equation x^2+y^2+Ax+By+C=0 where do A, B, and C come from

OpenStudy (zarkon):

that is if you multiply out \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

oh it was the diameter! good work

OpenStudy (zarkon):

yeah...took us a while but we figured that out ;)

OpenStudy (anonymous):

in the general equation x^2+y^2+Ax+By+C=0 where do A, B, and C come from??

OpenStudy (zarkon):

\[(x-h)^2+(y-k)^2=r^2\] \[x^2-2xh+h^2+y^2-2yk+k^2-r^2=0\] \[x^2+y^2+(-2h)x+(-2k)y+(h^2+k^2-r^2)=0\] \[A=-2h\] \[B=-2k\] \[C=h^2+k^2-r^2\]

OpenStudy (anonymous):

so you foiled out first. then completed the square?

OpenStudy (anonymous):

like how would you write (x-1/2)^2+(y)^2=1/4

OpenStudy (zarkon):

you don't need to complete the square

OpenStudy (zarkon):

\[(x-1/2)^2+(y)^2=1/4\] \[x^2-2\frac{1}{2}x+\frac{1}{4}+y^2-\frac{1}{4}=0\] \[x^2+y^2+\left(-2\frac{1}{2}\right)x+0y+0=0\]

OpenStudy (zarkon):

\[x^2+y^2-x=0\]